ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 150, 2005

Ball, K. Monotone factors of i.i.d. processes  205

Cao, W. See X. Wang

Chang, M-C. A sum-product estimate in algebraic division algebras 369
Hanke, T. A twisted Laurent series ring that is a noncrossed product 199

Junge, M., Z.-J. Ruan and D. Sherman A classification for 2-isometries of noncommutative L_{p}-spaces

285

Kalaj, D. On the Nitsche conjecture for harmonic mappings in ℝ^{2}

and ℝ^{3} 241

Li, L. See X. Wang

Martin, J. L. and V. Reiner Cyclotomic and simplicial matroids  229

Matet, P., C. Péan and S. Shelah Cofinality of normal ideals on

P_{κ}(λ) II 253

Péan, C. See P. Matet

Pfeifle, J. Long monotone paths on simple 4-polytopes  333

Reiner, V. See J. L. Martin

Ruan, Z.-J. See M. Junge

Sela, Z. Diophantine geometry over groups V_{1}: Quantifier elimination I

1

Shelah, S. See P. Matet

Sherman, D. See M. Junge

Solecki, S. Extending partial isometries  315

Wang, X., L. Li and W. Cao Discreteness criteria for Möbius groups acting on R̄^{n} 357